# Towards Universal Performance Modeling for Machine Learning Training on Multi-GPU Platforms

### v2

@@ III-C Communication Collective Performance Modeling
$t_{comm}=t_{cold\_start}+t_{per\_word}m_{size}.$ (1)
Distilled from a comprehensive study on multi-GPU communication on platforms by Li et al. [32] and a prior performance model work (LogCA [33]), we make a critical observation that eventually leads to the improvement of the above model:
@@
Therefore, we model the data movement latency as a piece-wise function of message size $m$:
@@
and $param$ include 4 parameters $L$, $x_{0}$, $k$, and $b$ that define the function
@@
This method only involves 8 parameters for each op (4 for sigmoid, 2 for region boundaries, 1 for startup latency, and 1 for maximum bandwidth). It is fast and simple for both curve fitting and prediction; it is also topology-agnostic and thus able to generalize to any communication pattern. We utilize the PARAM benchmark [34] and generate microbenchmark data for curve-fitting of both ops with equal message size starting from 4 bytes on each device and doubling it until exceeding the device memory. Test data is generated with random per-device message size for all-to-all and equal per-device message size for all-reduce. We calculate the input message size to the performance model as “the maximum of (the maximum of sent/received message size per device) across all devices”.
III-D Multi-GPU E2E Performance Modeling
As we mentioned, the main difference between single-GPU and multi-GPU performance modeling is that in multi-GPU scenarios, communication collectives like all-to-all and all-reduce simultaneously execute on multiple devices and trigger synchronizations and waiting across ranks and streams. This additional behavior significantly impacts runtime; this subsection describes how we simulate synchronizations and waiting during the execution analysis.
Before we dive deep into the multi-GPU E2E performance modeling algorithm, we need to first understand two types of synchronization in distributed training. They are:
@@ III-D Multi-GPU E2E Performance Modeling
These two types of synchronization are depicted in Figure 5. We expect that communication kernels’ launch time is different across ranks because of the different latency each rank takes to reach these ops due to unbalanced prior loads, such as data movement and computation. However, since communication kernels are synchronous, theoretically they must terminate at the same moment, while in reality they also terminate almost at the same time with negligible variance. This is the inter-rank synchronization, marked by the red dashed line in the plot. When it occurs, the time front of the communication stream on different ranks should be set to the same value during analysis. On the other hand, since the output data of the communication ops are used as inputs to certain successive ops, successive communication kernels, e.g., the last communication kernels on $S_{cm}$ on both ranks in Figure 5, should be launched at least after the first dependent kernel of the previous communication kernel (marked in blue dashed lines) rather than right after the previous communication kernel. This is called intra-rank/inter-stream synchronization. To reflect it, we set the communication stream time front to the same value as the launch time of the first dependent kernel.
@@
Figure 5: Two types of synchronization: inter-rank synchronization (red dashed line) and intra-rank/inter-stream synchronization (blue dotted line). For simplicity, we assume two GPUs and two streams ($S_{cp}$ and $S_{cm}$, for compute and communication respectively) per GPU, while CPU op calls are omitted in the plot. GPU kernels are represented by rectangles, and arrows indicate the data dependency between compute and communication kernels.
@@
We claim that both inter-rank and intra-rank synchronizations are the keys to accurately modeling ML workloads training performance on multi-GPU, or even more broadly, all types of workloads running on multi-heterogeneous-device platforms. This is because the best modeling method without them is to sum the kernel time per GPU stream and take the maximum as the E2E prediction result. It can possibly work for a few workloads that are completely dominated by one certain type of stream, e.g., GPT2 being dominated by the compute stream as shown above. However, in most cases, this will miss a large amount of GPU idle time (e.g., the gap between the red and blue lines in Figure 5) created by data dependencies and other possible waiting conditions during the execution, which will result in a significant and unacceptable underestimation of execution time, i.e., up to 60% as we show later in Section IV-B1.
@@
To take these synchronizations into consideration, we need the information on data dependency between ops in the ML model/workload, which is thankfully provided by the PyTorch execution graph (EG). We extend the critical-path-based algorithm proposed by Lin et al. [14] for single-GPU performance modeling to adapt to multi-GPU scenarios by incorporating inter-rank and intra-rank synchronizations, as shown in Algorithm 1. The new algorithm has a similar basic structure to the single-GPU version. Based on our observation, we assume two streams (one for compute and memory kernels and one for communication kernels) on each rank as that is the case for a vast majority of ML workloads. Instead of running one single process for training time prediction on a single GPU, the new algorithm runs $N$ processes in parallel, processing $N$ EGs and predicting execution time for $N$ ranks simultaneously (line 4). In addition to CPU time and GPU active time, we also track communication stream time $T_{cm}$ and compute/memory stream time $T_{cp}$ per process (line 6). At line 7, we also initialize a variable $last\_comm\_op$ to keep track of the dependencies between compute/memory ops and the last communication op. During the traversal of EGs, we detect intra-rank synchronization when a dependent op of the last communication op or a new communication op is encountered (lines 12–15); when the current op is a communication op, we do inter-rank synchronization after processing all kernels of it (lines 29–31) and update $last\_comm\_op$ for the next iteration (lines 36–38). At the end (lines 40–41), we calculate the local total time and GPU active time and do an all_gather operation across all ranks to collect their values, the maximums of which are returned as the predicted total time and GPU active time.
In our experiments to predict the multi-GPU training performance of DLRM models, we first generate 20 tasks with embedding tables randomly sampled from the DLRM open-source dataset. Among these 20 tasks, each of the 2021 and 2022 parts of the dataset contributes 5 heavy tasks (i.e., all sampled tables are heavy) and 5 normal tasks (i.e., tables can either be heavy or light). The total number of embedding tables per task is in the range of $(0.7\sim 1.3)\times\#\text{GPU}\times 13$ to resemble real-world workloads. To prevent out-of-memory errors, for each task we limit the total memory footprint of all embedding tables on each rank to be less than 80% of the GPU’s DRAM size. We set the embedding table sharder to size_lookup_greedy based, i.e., the cost of each table is estimated as $L\times D\times\log(E)$. For each training iteration in the E2E test of each task, we randomly sample a mini-batch of data (with batch size set to 512/1024/2048/4096) corresponding to each table selected in the task from the dataset and distribute it to each rank based on the sharding scheme. The overhead statistics (mean latency, etc) of PyTorch ops are aggregated from all collected traces and shared by all tested workloads. The actual/predicted time is measured/calculated by averaging over 30 iterations. It takes about a day to run the benchmark and analysis for all 80 (20 workloads $\times$ 4 batch sizes) tasks, while it takes less than 20 minutes to predict their performance.
@@
1:Input: $EG_{0,1,\ldots N-1}$ of an ML workload trained on a single-node N-GPU platform, one EG per rank; Kernel performance models $\{M\}$; Overheads $Ov$.
@@
41:     Synchronize $T$ and $gpu\_time$ across all processes and take their maximums.
@@
III-E Embedding Lookup Kernel Modeling
The performance modeling of Embedding lookup (EL), an indispensable and commonly dominating [14] operator in many recommendation models including DLRM, is challenging in a few aspects. First, the dimensions and parameters of every single embedding table, specifically the number of embeddings ($E$), the embedding dimension ($D$), and the pooling factor ($L$) can all be different in real workloads and data. Second, the distribution of input data to each table varies considerably across each batch. Both these two factors significantly increase the difficulty of applying a heuristic-based performance model to the EL op, since it is almost impossible to accurately estimate the L2 cache hit rate and data movement and thus predict the latency of the op. Previous work such as Lin et al. [14] only considers modeling the performance of the case when $D$, $E$, and $L$ are all fixed for each table and data distribution is uniform. Therefore their solution is not sufficient to deal with real-world workloads and data.
@@ III-E Embedding Lookup Kernel Modeling
Figure 6: Generation of reuse factors (RF) from embedding lookup indices. (a) is an example of lookup indices of a batch of 4 samples on one embedding table. In (b), the x-axis is the indices and the y-axis is the count of these indices. In (c), the x-axis is the count of accesses and the y-axis is the number of such indices.
@@
To address this issue, we adopt a straightforward and clear expression of describing EL’s input data distribution using reuse factors (RF), as introduced by Meta’s open-source DLRM dataset [29]. Figure 6 shows an example of generating reuse factors (RF) from embedding lookup indices. First, the histogram of per-table index counts in a batch of input data is calculated ((a) to (b)). Then, the “histogram of histogram”, i.e., how many indices are accessed once, twice, thrice, etc., is calculated, which is thus the RF of this batch on a certain table ((b) to (c)). In practice, the bins for this step have sizes of exponents of 2, e.g., $[0,1)$, $[1,2)$, $[2,4)$, $[4,8)$, etc. Finally, the counts of every bin are normalized into the range of $[0,1]$. In this way, the distribution of lookup indices is described as by what probability any row of a table is hit by a certain range of times, such as $2^{m}$ to $2^{n}$ times where $m$ and $n$ are both integers and $m<n$. Meta’s DLRM dataset provides the RF of 17 bins for each data file, and we follow this convention in our experiments. The advantages of RF are its low overhead in terms of both processing time and parameter size. The calculation is relatively simple, and thus it is trivial to obtain RF values of each batch of input data on the fly before the training of a batch is actually started. Because each table has 17 RF values and the number of tables residing on each rank is usually limited, e.g., around 10, the total number of RF values for each batch of input data remains acceptable.
@@
Figure 7: RF values of the top 10 most visited tables among 6000 random samples from the DLRM open-source dataset for the EL microbenchmark. Red dashed lines represent the cumulative distribution function (CDF) of the overall RF values of each table in the whole dataset, while each thin gray line represents the CDF of RF values of a sample that visits a certain table. It can be clearly observed that the data distribution (represented by RF values) of each data batch is different and sometimes distant from that of the whole dataset. Compared to overall RF values, the RF values of a data batch tend to concentrate towards 0 along the X axis.
@@
We adapt the cost model proposed by Zha et al. [28] for our use of real-time performance modeling of EL. Zha et al. create a multi-head MLP cost model to simultaneously predict the cost of the forward time, backward time, and communication time of a sharding scheme of EL as a whole on multiple devices. Our method is different from theirs in two aspects. 1) We handle these three constituent times separately in the granularity of ops to accommodate our robust E2E algorithm shown in Section III-D. This better simulates and models the actual execution since considering these three times as a whole misses the opportunity to explain the computation-communication overlaps (also described in Section III-D). 2) Instead of using RF values of each table, we use RF values of each batch for better prediction based on the reason stated in the last paragraph. We construct our MLP performance model resembling one of the heads in Zha et al.’s cost model and match those used for other kernels for convenient training and inference. To enable EL’s execution incorporating flexible $D$, $E$, and $L$, in this work, we use a popular, efficient, and flexible batched-EL implementation provided by FBGEMM [35], an open-source high-performance kernel library for training and inference on both CPUs and GPUs. We first create a microbenchmark dataset by randomly sampling tables and batches (e.g., 10 tables out of 856, and a batch of 1024 lookups out of 65536) from the DLRM open-source dataset, and make sure that the size of the sampled tables in each sample will not exceed the DRAM memory size of the selected GPU. Note that while in the previous work, the cost model is trained with RF values of the whole dataset, it is improper to do so for real-time performance modeling because RF values tend to spread in a larger range. This fact is easy to understand intuitively. Suppose the batch size is small, the number of times each row in a table is hit tends to concentrate close to 0 because there are few indices in the batch. If we consider the whole dataset as a huge batch (e.g., size 65536 for the DLRM dataset we use), it is very likely that there are some rows in a table to be hit many times, such as 500 times, which could never appear with a small batch size such as 256. In a word, these two distributions are different (as shown in Figure 7), and thus the distribution of the whole dataset is not representative of a batch of data sampled from it. Therefore, we calculate the RF values of each batch and store them together with the execution time as part of the training data. Finally, we train and validate the performance models with the microbenchmark results as the input dataset that is split into a training set and a test set with a split factor of 0.8.
III-F Additional Ops Performance Modeling Support
@@
To support Transformer-based NLP models, We also add kernel performance models for additional ops including layer_norm, dropout, and element-wise ops such as gelu and tanh. Using the same approach as the base model[14], we predict the latency of layer_norm and dropout with ML-based models trained with microbenchmark data of PyTorch ops and predict element-wise ops using the roofline model.
IV Evaluation and Analysis
@@
We evaluate our kernel and E2E performance models with PyTorch v2.0 along with FBGEMM v0.4.1, CUDA 11.7, and Python 3.9. The performance model of FBGEMM’s embedding lookup kernel is evaluated on single NVIDIA GV100 and A100 (40 GB) GPUs, while that of both communication kernels and multi-GPU E2E are evaluated on two multi-GPU platforms, including 4xGV100 equipped with 48-core Intel(R) Xeon(R) Gold 6146 CPU @ 3.20 GHz, and 4xA100 equipped with GCP’s a2-highgpu-4g with 48-vCPU. The GPU communication topologies of these two platforms are shown in Figure 8. We use data sampled from the DLRM open-source dataset [29] as both microbenchmark data for FBGEMM embedding lookup kernel performance model training and verification, and input data for embedding lookup of the DLRM models in E2E tests. Multiple pieces (.pt files) of the dataset are merged for later use. The multi-GPU E2E evaluation in this work covers DLRM model training (code adapted from https://github.com/facebookresearch/dlrm) and fine-tuning of natural language processing (NLP) models such as BERT [15], GPT2 [16], and XLNet [17] (all model implementations called from HuggingFace’s Transformers library [36]). We use PyTorch’s DistributedDataParallel (DDP) in distributed training of ML workloads in our experiments for no GIL contention and less model replication and data movement overheads [37]. In addition, we manually insert a barrier at the beginning of each batch for better profiling of communication ops; we set the bucket size of gradient bucketing [38] to the default value (25) for all our experiments, as we observe that varying it only causes negligible change of performance.
@@
IV-A Kernel Performance Modeling
We obtain less than 10% GMAE and MAPE prediction errors for all kernel performance models shown in Table I. Particularly, the adjustment of all-to-all message size yields a low prediction error of latency, implying that the operation is bounded by the biggest per-device data bulk sent from or received by one certain device. Specifically, we present the fitted curves for all-to-all benchmark data on both platforms in Figure 9. We observe that the practical problem size for all-to-all in DLRM workloads lies in section 2 (the transitional section), which justifies our improvement on Equation 1.
@@ IV-A Kernel Performance Modeling
IV-B Multi-GPU E2E Performance Modeling
IV-B1 DLRM on Multi-GPU Platforms
We present the statistics of all E2E performance modeling tests on the two multi-GPU platforms in Table II, and the prediction error and the reference time of each task in Figure 10. The baseline result to be compared with our prediction in Figure 10 is given by the maximum sum of kernel active time of each GPU stream. We can see that this baseline prediction, yielding error values higher than 60%, is not sufficient to be used as the predicted E2E time per iteration because no idle time nor waiting time caused by data dependency between the streams is considered. Instead, our enhanced algorithm can accurately predict both normal and heavy tasks with high accuracy, with an overall geomean prediction error of 5.21%. The majority of prediction results on both devices underestimate the actual time. One possible reason is that with big batch sizes, the communication time dominates the per-iteration time so the syncing and waiting time among all ranks also increases and contributes to the per-iteration time. The remaining tests that overestimate might be explained by the overestimation of GPU idle time caused by CPU overheads when the workload is latency-bound. Device-wise, the behavior of prediction errors, such as geomean/minimum/maximum error and trends when the batch size changes, does not deviate much, which justifies the consistency and stability of our prediction algorithm across platforms.

### v3

@@ IV-A Communication Collective Performance Modeling
$t_{comm}=t_{cold\_start}+t_{per\_word}m_{size}$ (1)
with $t_{cold\_start}$ representing the fixed overhead of data transaction, and $t_{per\_word}$ being the per-word data transaction time of the communication medium. Distilled from a comprehensive study on multi-GPU communication on platforms by Li et al. [32] and a prior performance model work (LogCA [33]), we make a critical observation that eventually leads to the improvement of the above model:
@@
Therefore, we model the data movement latency as a piece-wise function of message size $m$ with $m_{1}$ and $m_{2}$ being boundary sizes separating the three regions:
@@
and $param$ include 4 parameters $L$, $x_{0}$, $k$, and $b$ that define a standard sigmoid function:
@@
This method only involves 8 parameters for each op (4 for sigmoid, 2 for region boundaries, 1 for startup latency, and 1 for maximum bandwidth). It is fast and simple for both curve fitting and prediction; it is also topology-agnostic and thus able to generalize to any communication pattern. We utilize the PARAM benchmark [34] and generate microbenchmark data to fit curves for both ops with equal message size starting from 4 bytes on each device and doubling it until exceeding the device memory. Test data is generated with random per-device message size for all-to-all and equal per-device message size for all-reduce. We calculate the input message size to the performance model as “the maximum of (the maximum of sent/received message size per device) across all devices”.
IV-B Multi-GPU E2E Performance Modeling
As we mentioned in Section III-B, the main difference between single-GPU and multi-GPU performance modeling is that in multi-GPU scenarios, communication collectives like all-to-all and all-reduce simultaneously execute on multiple devices and trigger synchronizations and waiting across ranks and streams. This additional behavior significantly impacts runtime. This subsection describes how we simulate synchronizations and waiting during the execution analysis.
Before we dive deep into the multi-GPU E2E performance modeling algorithm, we need to first understand two types of synchronization in distributed training (depicted in Figure 5):
@@ IV-B Multi-GPU E2E Performance Modeling
We expect communication kernels’ launch time to differ across ranks because each rank’s latency to reach these ops due to unbalanced prior loads, such as data movement and computation, is different. However, since communication kernels are synchronous, theoretically they must terminate simultaneously. In reality, they do terminate almost simultaneously with negligible variance. This is the inter-rank synchronization, marked by the red dashed line in Figure 5. When it occurs, the time fronts of the communication streams on different ranks should be set to the same during analysis. Nevertheless, since the output data of the communication ops are used as inputs to certain successive ops, successive communication kernels, e.g., the last communication kernels on $S_{cm}$ on both ranks in Figure 5, should be launched at least after the first dependent kernel of the previous communication kernel (marked in blue dashed lines) rather than right after the previous communication kernel. This is called intra-rank/inter-stream synchronization. To reflect it, we set the communication stream time front to the same as the launch time of the first dependent kernel.
@@
Figure 5: Inter-rank synchronization (red dashed line) and intra-rank/inter-stream synchronization (blue dotted line). For simplicity, we assume two GPUs and two streams ($S_{cp}$ and $S_{cm}$, for compute and communication respectively) per GPU, while CPU op calls are omitted in the plot. Rectangles represent GPU kernels, and arrows indicate the data dependency between compute and communication kernels.
@@
We claim that the modeling of both inter-rank and intra-rank synchronizations are the keys to accurately modeling ML workloads training performance on multi-GPU, or even more broadly, all types of workloads running on multi-heterogeneous-device platforms. This is because the best modeling method without considering them is to sum the kernel time per GPU stream and take the maximum as the E2E prediction result. It might work for a few workloads dominated by one certain type of stream, e.g., GPT2 being dominated by the compute stream as shown above. However, in most cases, this will miss a large amount of GPU idle time (e.g., the gap between the red and blue lines in Figure 5) created by data dependencies and other possible waiting conditions during the execution, leading to a significant and unacceptable underestimation of execution time, i.e., up to 60% as we show later in Section V-B1.
@@
To consider these synchronizations, we need the information on data dependency between ops in the ML model/workload, which is thankfully provided by the PyTorch execution trace (ET). We extend the critical-path-based algorithm proposed by Lin et al. [14] for single-GPU performance modeling to adapt to multi-GPU scenarios by incorporating inter-rank and intra-rank synchronizations, as shown in Algorithm 1. The new algorithm has a similar basic structure to the single-GPU version. Based on our observation, we assume two streams (one for compute and memory kernels and one for communication kernels) on each rank as that is the case for most ML workloads. Instead of running a single process for training time prediction on a single GPU, the new algorithm runs $N$ processes in parallel, processing $N$ EGs and predicting execution time for $N$ ranks simultaneously (line 4). In addition to CPU time and GPU active time, we also track communication stream time $T_{cm}$ and compute/memory stream time $T_{cp}$ per process (line 6). At line 7, we also initialize a variable $last\_comm\_op$ to keep track of the dependency between compute/memory ops and the last communication op. To recap, T1–T5 in line 9 represents the five types of execution overhead on the CPU of each op, which are summarized in our previous work [14]. During the traversal of ETs, we detect intra-rank synchronization when an op that has data dependency on the last communication op or a new communication op is encountered (lines 12–15); when the current op is a communication op, we do inter-rank synchronization after processing all kernels of it (lines 29–31) and update $last\_comm\_op$ for the next iteration (lines 36–38). At the end (lines 40–41), we calculate the local total time and GPU active time and do an all_gather operation across all ranks to collect their values, the maximums of which are returned as the predicted total time and GPU active time.
@@
1:Input: $ET_{0,1,\ldots N-1}$ of an ML workload trained on a single-node N-GPU platform, one execution trace per rank; Kernel performance models $\{M\}$; Overhead statistics $Ov$.
@@
41:     Synchronize $T$ and $gpu\_time$ across all processes and return their maximums.
@@
IV-C Embedding Lookup Kernel Modeling
The performance modeling of EL, an indispensable and commonly dominating [14] operator in many recommendation models including DLRM, is challenging in a few aspects. First, the dimensions and parameters of every single embedding table, specifically the number of embeddings ($E$), the embedding dimension ($D$), and the pooling factor ($L$) can all be different in real workloads and data. Second, the input data distribution of each table varies considerably across each batch. Both factors significantly increase the difficulty of applying a heuristic-based performance model to the EL op, since it is almost impossible to estimate the L2 cache hit rate and data movement accurately and thus predict the latency of the op. Previous work such as Lin et al. [14] only considers modeling the performance of the case when $D$, $E$, and $L$ are all fixed for each table with uniform input data distribution, making their solution insufficient to deal with real-world workloads and data.
@@ IV-C Embedding Lookup Kernel Modeling
Figure 6: Generation of reuse factors (RF) from EL indices. (a) is an example of lookup indices of a batch of 4 samples on one embedding table. In (b), the x-axis is the indices and the y-axis is the count of these indices. In (c), the x-axis is the count of accesses and the y-axis is the number of such indices.
@@
To address this issue, we adopt a straightforward and clear expression of describing EL’s input data distribution using reuse factors (RF), as introduced by Meta’s open-source DLRM dataset [29]. Figure 6 shows an example of generating reuse factors (RF) from EL indices. First, the histogram of per-table index counts in a batch of input data is calculated ((a) to (b)). Then, the “histogram of histogram”, i.e., how many indices are accessed once, twice, thrice, etc., is calculated, which is thus the RF of this batch on a certain table ((b) to (c)). In practice, the bins for this step have sizes of exponents of 2, e.g., $[0,1)$, $[1,2)$, $[2,4)$, $[4,8)$, etc. Finally, the counts of every bin are normalized into the range of $[0,1]$. In this way, the distribution of lookup indices is described as by what probability any row of a table is hit by a certain range of times, such as $2^{m}$ to $2^{n}$ times where $m$ and $n$ are both integers and $m<n$. Meta’s DLRM dataset provides the RF of 17 bins for each data file, and we follow this convention in our experiments. The advantages of RF are its low overhead in terms of processing time and parameter size. The calculation is relatively simple, and thus it is trivial to obtain RF values of each batch of input data on the fly before the training of a batch is started. Because each table has 17 RF values and the number of tables residing on each rank is usually limited, e.g., around 10, the total number of RF values for each batch of input data remains acceptable.
@@
Figure 7: RF values of the top 10 most visited tables among 6000 random samples from the DLRM open-source dataset for the EL microbenchmark. Red dashed lines represent the cumulative distribution function (CDF) of the overall RF values of each table in the whole dataset, while each thin gray line represents the CDF of RF values of a sample that visits a certain table. It can be observed that the data distribution (represented by RF values) of each data batch is different and sometimes distant from that of the whole dataset. Compared to overall RF values, the RF values of a data batch tend to concentrate towards 0 along the X axis.
@@
We adapt the cost model proposed by Zha et al. [28] and use it in real-time performance modeling of EL. Zha et al. create a multi-head MLP cost model to simultaneously predict the cost of the forward time, backward time, and communication time of a sharding scheme of EL as a whole on multiple devices. Our method is different in two aspects. 1) We handle these three constituent times separately in the granularity of ops to accommodate our robust E2E algorithm shown in Section IV-B. This better simulates and models the actual execution since considering these three times as a whole misses the opportunity to explain the computation-communication overlaps (described in Section IV-B). 2) Instead of using RF values of each table, we use RF values of each batch for better prediction based on the reason stated in the last paragraph. We construct our MLP performance model resembling one of the heads in Zha et al.’s cost model and match those used for other kernels for convenient training and inference. To enable EL’s execution incorporating flexible $D$, $E$, and $L$, we use a popular, efficient, and flexible batched-EL implementation provided by FBGEMM [35], an open-source high-performance kernel library for training and inference on both CPUs and GPUs. We first create a microbenchmark dataset by randomly sampling tables and batches (e.g., 10 tables out of 856, and a batch of 1024 lookups out of 65536) from the DLRM open-source dataset, and make sure that the size of the sampled tables in each sample will not exceed the DRAM memory size of the selected GPU. Although Zha et al. trained the cost model with RF values of the whole dataset, it is improper to do so for real-time performance modeling because RF values tend to spread in a larger range. This fact is easy to understand intuitively. Suppose the batch size is small, the number of times each row in a table is hit tends to concentrate close to 0 because there are few indices in the batch. If we consider the whole dataset as a huge batch (e.g., size 65536 for the DLRM dataset we use), there are likely some rows in a table to be hit many times, such as 500 times, which could never appear with a small batch size such as 256. In a word, these two distributions are different (as shown in Figure 7), and thus the distribution of the whole dataset is not representative of a batch of data sampled from it. Therefore, we calculate the RF values of each batch and store them together with the execution time as part of the training data. Finally, we randomly split the microbenchmark result dataset into a training set and a test set with a split factor of 0.8, and use them to train and validate the performance models.
IV-D Additional Ops Performance Modeling Support
@@
To support Transformer-based NLP models, We also add kernel performance models for additional ops including layer_norm, dropout, and element-wise ops such as gelu and tanh. By applying principles mentioned in our previous work [14], ML-based models suit layer_norm and dropout for their hidden kernel implementation detail in PyTorch, while the roofline model suits element-wise ops for their simplicity. Therefore we predict the latency of layer_norm, dropout with ML-based models trained with microbenchmark data of PyTorch ops for and predict gelu and tanh using the roofline model.
V Evaluation and Analysis
@@
We evaluate our kernel and E2E performance models with PyTorch v2.0, FBGEMM v0.4.1, CUDA 11.7, and Python 3.9. The performance model of FBGEMM’s embedding lookup kernel is evaluated on single NVIDIA GV100 and A100 (40 GB) GPUs, while that of both communication kernels and multi-GPU E2E are assessed on two multi-GPU platforms, including 4xGV100 equipped with 48-core Intel(R) Xeon(R) Gold 6146 CPU @ 3.20 GHz, and 4xA100 equipped with GCP’s a2-highgpu-4g with 48-vCPU. Figure 8 shows the GPU communication topologies of these two platforms. We use data sampled from the DLRM open-source dataset [29] as both microbenchmark data for FBGEMM embedding lookup kernel performance model training and verification, and input data for embedding lookup of the DLRM models in E2E tests. Multiple pieces (.pt files) of the dataset are merged for later use. The multi-GPU E2E evaluation in this work covers DLRM model training (code adapted from https://github.com/facebookresearch/dlrm) and fine-tuning of natural language processing (NLP) models such as BERT [15], GPT2 [16], and XLNet [17] (all model implementations called from HuggingFace’s Transformers library [36]). We use PyTorch’s DistributedDataParallel (DDP) in distributed training of ML workloads in our experiments for no GIL contention and less model replication and data movement overheads [37]. In addition, we manually insert a barrier at the beginning of each batch for better profiling of communication ops; we set the bucket size of gradient bucketing [38] to the default value (25) for all our experiments, as we observe that varying it only causes a negligible performance change.
@@
V-A Kernel Performance Modeling
We provide prediction errors in two metrics: GMAE (geometric mean absolute error) to reduce the impact of outliers that is common when measuring E2E latency, and MAPE (mean absolute percentage error) as an intuitive measurement of deviation as an (always positive) percentage. We obtain less than 10% both GMAE and MAPE for all kernel performance models shown in Table I. Particularly, the adjustment of all-to-all message size yields a low prediction error of latency, implying that the operation is bounded by the biggest per-device data bulk sent from or received by one certain device. Specifically, we present the fitted curves for all-to-all benchmark data on both platforms in Figure 9. We observe that the practical problem size for all-to-all in DLRM workloads lies in section 2 (the transitional section), which justifies our improvement on Equation 1.